OpenStudy (anonymous):

x + y = 12 x - y = 10 What is the value of the x-determinant for the system shown? a)-22 b)-12 c)-2

OpenStudy (anonymous):

-22. I think you are trying to find Dx, so you'd set up the determinant as |12 1 | |10 -1| =-12-10 = -22